# AI-Driven Control of Chaos: A Transformer-Based Approach for Dynamical Systems

### v1

@@ IV Results
Figure 4 illustrates the neural network’s accuracy by comparing the true safety function of a random map with the safety functions predicted by the trained models. The neural network predicts the safety function value for a single initial condition at a time. To reconstruct the complete safety function, $1000$ initial conditions were sampled from the region $Q$, and the same time series data was used as input for each prediction. Networks trained on a larger number of points produce predictions that closely approximate the true safety function, though they tend to slightly underestimate the curve. Conversely, models trained on fewer points are more likely to overestimate the safety function.
In addition to the predicted safety functions, each plot includes a cyan curve representing the maximum control across all initial conditions in $Q$ over $1000$ iterations following the controls suggested by the predicted safety function. This maximum control value is by definition, the value of the safety function at that point, and therefore is computed according to Eq. 1, iterating the safety function over this interval and initializing $U_{0}$ with the predicted safety function. Notably, the maximum control typically aligns closely with the true safety function. However, in regions of minimum control, a lower bound emerges where all orbits exhibit the same maximum control value.
Figure 5 illustrates the observed behavior in an orbit with initial condition $x_{n}=0.7$ iterated $1000$ times. The safe set is derived from the safety function, where the value of the safety function at the initial condition determines the minimum control needed to keep the orbit controlled indefinitely. The safe set includes all regions in $Q$ where the safety function is less or equal to this value.
@@
Using the true safety function in Fig. 5a), the safe set identifies regions where the orbit remains stable indefinitely under the suggested control. In Fig. 5b), the predicted safety function, derived from a model trained with $2000$ points, introduces errors in the safe set. Overestimation of the safety function leads to underestimation of the safe set (red regions), while underestimation of the safety function causes overestimation of the safe set (orange regions). A zoomed in view near $x_{n}=0.7$, shown in the top right corner of Fig. 5b), highlights these discrepancies.
@@
Figure 4: Comparison of predicted safety functions with the true safety function across various models. Each subplot illustrates the similarity of the predictions with their ground truths. As the number of sampled points increases, the predicted safety function aligns more closely with the true safety function. The thin blue line denotes the maximum control for each initial condition over $1000$ iterations in the map using the control values suggested by the predicted safety function, by iterating the safety function using Eq. 1 but initializing $U_{0}$ with the values of the predicted safety function, we can measure the maximum control value for each orbit.
A significant challenge in this model free approach lies in accurately capturing the lower control regions of the true safety function, particularly when data is sparse. In such scenarios, the estimated safety function often diverges from the true safety function in these regions. This discrepancy arises because sharp variations in the lower control zones depend on data points that are frequently absent during sampling, hindering precise detection and representation. This phenomenon aligns with previous studies on dynamical systems subject to minimal disturbance, which demonstrate that as control decreases, the safe set of orbits within a region $Q$ forms a Cantor-like structure, eventually vanishing from the system capeans2019new .
Regarding the performance of the trained models, Fig. 6 illustrates the evolution of the mean squared error (MSE) in estimating safety functions as the noise value $\xi$ in the source maps increases. This analysis involved the computation of $1500$ safety functions derived from random maps. Specifically, for $100$ values of $\xi$, ranging from $0.001$ to $0.1$, $15$ safety functions were generated using maps with the same $\xi$. For each safety function, the mean MSE was calculated in two steps: first, determining the average MSE across all $1000$ points defining the curve; and second, averaging these MSE values over the $15$ safety functions.
@@
The results demonstrate that the models estimate safety functions more accurately for maps with lower noise upper bounds, with performance deteriorating as the noise value increases. Nevertheless, the predictions remain reliable across the entire range of noise levels. This trend aligns with a training induced bias, as described in Appendix B, where systems with lower noise were prioritized to enhance the estimation of the lower control regions of the safety function.
To conclude the analysis of the aforementioned dataset, Table 1 provides a detailed summary of each training model’s performance, including the mean prediction time per safety function with $1000$ points and its standard deviation, as well as the mean MSE of the dataset and its standard deviation.
@@
Table 1: Model performance summary: Increasing the number of sampled points generally improves accuracy (lower MSE) but extends computation time. Models with $1000$ or $500$ points achieve similar MSE to the $2000$ point model under evenly distributed noise. Additionally, models with fewer than $500$ points show minimal computation time variation, likely due to hardware constraints discussed in Section III. This table, in conjunction with Fig. 6, highlights the trade-off between model accuracy, prediction time, and the effects of model size on prediction error across different noise levels.
As expected, increasing the amount of information sampled from a system, results in longer prediction times and reduced MSE. However, for models processing $500$ points or fewer, no significant variation in computation time was observed. This consistency in computation time is likely attributed to the hardware specifications used for executing the models, as discussed in Sec. III. This lower threshold reflects the computational overhead of the programs that can be optimized to meet a application requirement.
@@
V Conclusion
In this study, we propose a transformer-based framework designed to predict the safety functions of dynamical systems without the need for an explicit physical model. This model free approach offers notable advantages over traditional algorithms. Specifically, our framework achieves competitive accuracy while drastically reducing computational demands, particularly in scenarios where classical methods are computationally prohibitive.
@@ V Conclusion
A key strength of the transformer-based approach is its significantly lower computational complexity compared to classical methods. The classical computation of safety functions, as defined in Appendix A, involves an iterative process with a computational complexity of $\mathcal{O}(K\cdot M\cdot N\cdot 2^{d})$, where $K$ is the number of iterations required for the safety function to converge, $M$ is all the possible disturbance scenarios affecting a point, $N$ is the number of neighbors per point, and $d$ is the dimensionality of the system. This iterative nature makes classical methods computationally prohibitive for higher dimensional systems, especially as the number of neighbors or required iterations increases.
@@
We also explored the trade-off between data volume, predictive accuracy, and computation time for safety functions composing of $1000$ points. As shown in Table 1, increasing sample size generally improves accuracy and robustness but at the cost of increased training time and computational demands. For example, the $1000$ point model achieved an MSE that was only $3.6\%$ higher than the $2000$ point model, yet it performed predictions $2.84$ times faster. These results suggest that selecting an optimal balance between sampled points and computational cost can yield efficient and accurate performance tailored to specific applications. Further research could explore strategies such as adaptive sampling and padding to enhance model performance with limited or imbalanced data.
@@ Appendix A Classical computation of safety functions
This appendix provides a detailed methodology for the classical computation of safety functions. These are a critical tool in extending the partial control method, enabling the precise determination of minimal control bounds necessary to sustain trajectories within a predefined region $Q$ of the phase space.
@@ A.1 Computation with disturbances
Figure 8: Scheme of a map affected by a bounded disturbance distribution. In this case, to compute the function $U_{k+1}[q_{i}]$, we have to consider all disturbed images $f(q_{i},\xi_{s})$ corresponding to the point $q_{i}$. Then compute all the corresponding control bounds, and finally extract the maximum control among them all.
@@
Comments: -Note that the $u[q_{i},\xi_{s},q_{j}]$ values remain unchanged every iteration of the while loop so compute them once and save them
A.2 Safety function and safe sets
The safety function $U_{\infty}(q)$ directly determines the safe set for a given control bound $u$. Specifically:
Safe $\text{Safe Set}=\{q\in Q:U_{\infty}(q)\leq u\}.$ (8)
This safe set determines the points in $Q$ that can be controlled within the specified limit, enabling the development of tailored interventions to sustain the system in a desired state.
A.3 Calculation of safety functions
@@
The true safety function is then computed iteratively using Eq. 7, which determines the minimum control required for each initial condition $q_{i}\in Q$ to ensure the controlled orbit remains within $Q$ indefinitely. During each iteration, the algorithm evaluates all potential noisy outputs and calculates the control necessary to keep the system within the safe set for every point in $Q$.
@@ B.1 Calculation of random dynamical systems
The number of functions to be used $M$ is randomly determined from predefined ranges for each set: $4$ to $8$ from Set A, $2$ to $4$ from Set B, $5$ to $8$ from Set C, and $1$ to $2$ from Set D. These ranges are designed to ensure diversity in the generated functions. Once the random functions $g_{i}(x)$ are selected from the respective sets, a weight $c_{i}$ is assigned to each function. These weights are random coefficients uniformly sampled from the range $[0,1]$. The weighted selected functions are then combined into a single expression $f(x)$ as shown in Eq.  9.
$f(x)=\sum_{i=1}^{M}c_{i}g_{i}(x),$ (9)
@@ B.2 Time series simulation
Regarding the system orbits, we iterated $1000$ random initial conditions under the system’s dynamics until they either exited region $Q$ or completed $50$ iterations. The time series evolved according to $x_{n+1}=f(x_{n})+\epsilon_{n}$. If an orbit exits region $Q$, it is truncated and padded with $-1$ values to achieve a consistent length of $50$ points. This process generates a time series with a shape $(1000,50)$ for each random function.
@@
Set A Set B Set C Set D
$x^{2}(1-x)^{2}$ $\lvert(1-x^{2})\sin^{2}(\pi x)\rvert$ $\lvert(1-e^{-5x})e^{-5(1-x)}x(1-x)\rvert$ $\lvert\sin(2\pi x)x(1-x)\rvert$
$\sin^{2}(\pi x)x(1-x)$ $\lvert(x-x^{2})\cos^{2}(\pi x)\rvert$ $\lvert\log(1+9x)(1-x)^{2}\rvert$ $\lvert x(1-x)(\text{mod}(10x,1)>0.5)\rvert$
$(4+r)x(1-x)^{2}$ $\lvert x(1-x)(1+x-x^{2})\rvert$ $\lvert(x-x^{2})^{2}\sin(2\pi x)\rvert$ $(3+r)\lvert\text{mod}(10x,1)-0.5\rvert x(1-x)$
$\sqrt{x}(1-x)^{2}$ $\lvert(1-x)\sin(\pi x)x\rvert$ $\lvert x^{2}(1-x)\cos^{2}(\pi x)\rvert$ $\lvert\lfloor 10x-5\rfloor(1-x)x\rvert$
$x^{3}(1-x)^{3}$ $\lvert(x(1-x))^{2}\rvert$ $\lvert x(1-x^{4})\rvert$ $(4+r)\lvert\cos(\pi x)\rvert x(1-x)$
$(1-\cos(\pi x))x(1-x)$ $\lvert x^{3}(1-x)(2-x)\rvert$ $\lvert e^{-2(x-0.5)^{2}}x(1-x)\rvert$ $\lvert\lceil 10x-5\rceil(1-x)x\rvert$
$\tanh^{2}(4x-2)x(1-x)$ $\lvert x(1-x)e^{-x}\rvert$ $\lvert(x(1-x))^{3}\rvert$ $\lvert(\text{mod}(20x,1)>0.5)(1-x)x\rvert$
$(1-e^{-4x})^{2}x(1-x)$ $\lvert(1-x)(1-\cos(2\pi x))\rvert$ $\lvert(\sin(2\pi x)+1)x(1-x)\rvert$ $\lvert\tan(\pi(x-0.5))(1-x)x\rvert$
$(2+r)\min\left(\frac{x}{c},\frac{1-x}{1-c}\right)^{2}x(1-x)$ $\lvert\sin^{2}(\pi x)x^{2}(1-x)\rvert$ $\lvert(x\log(x+10^{-5}))(1-x)\rvert$ $\lvert\text{mod}(\lfloor 5x\rfloor,2)(1-x)x\rvert$
$\lvert(1-e^{-x})(1-x)x\rvert$ $\lvert\sqrt{x}(1-x^{3})\rvert$ $\lvert\text{sawtooth}(2\pi x)(1-x)x\rvert$
$\lvert\cos(\pi x)x^{2}(1-x)\rvert$ $\lvert x^{3}(1-x^{5})\rvert$ $\lvert\text{sign}(\cos(2\pi x))(1-x)x\rvert$
$\lvert(x^{2}-x)\cos(\pi x)\rvert$ $\lvert\text{mod}(50x,1)(1-x)x\rvert$
$\lvert(x(1-x))\sin^{2}(4\pi x)\rvert$ $\lvert\cos(4\pi x)(1-x)x\rvert$
$\lvert(x(1-x))\cos^{2}(6\pi x)\rvert$
$\lvert(x(1-x))(\sin^{2}(2\pi x)+\cos^{2}(2\pi x))\rvert$
$\lvert(x(1-x))e^{-10(x-0.5)^{2}}\rvert$
$\lvert(x(1-x))(1+\cos(10\pi x))\rvert$
$\lvert(x(1-x))\lvert\sin(8\pi x)\rvert\rvert$
$\lvert(x(1-x))(1-\cos^{2}(12\pi x))\rvert$
$\lvert(x(1-x))(1+\sin^{2}(6\pi x))\rvert$
$\lvert(x(1-x))(1+e^{-5(x-0.3)^{2}}+e^{-5(x-0.7)^{2}})\rvert$
$\lvert(x(1-x))(1+\sin(4\pi x)\cos(4\pi x))\rvert$
Table 2: Table of functions for Sets A, B, C, and D. Set A functions are inspired by well known chaotic maps and include logistic, sinusoidal, and polynomial components. Set B functions are smooth and designed to represent oscillatory and exponential behaviors. Set C functions add complexity by incorporating logarithmic and Gaussian components. Set D functions introduce sharp transitions and discontinuities. The algebraic variables denoted by $r$ and $c$ are random uniform variables between 00 and $1$. Regarding the non obvious functions, log represents the base $10$ logarithm, Sawtooth represent a saw wave, Sign $f(x)$ returns the sign of the output value the function, $\lfloor f(x)\rfloor$, and $\lceil f(x)\rceil$ denote floor and ceiling over a function respectively, and finally mod$(a,b)$ is the modulo operation that returns the remainder of the division of $a/b$.

### v2

@@ 4 Results
Figure 4 illustrates the neural network’s accuracy by comparing the true safety function of a map generated randomly with the safety functions predicted by the trained models. The neural network predicts the value of the safety function $U_{\infty}(x_{k})$ for a single input state $x_{k}$. To reconstruct the complete safety function, $1000$ initial conditions were sampled from the region $Q$, and the same time series data was used as input for each prediction. Networks trained on a larger number of points produce predictions that closely approximate the true safety function, though they tend to slightly underestimate the curve. Conversely, models trained on fewer points are more likely to overestimate the safety function.
In addition to the predicted safety functions, each plot includes a cyan curve representing the maximum control across all initial conditions in $Q$ over $1000$ iterations following the controls computed with Eq. 4 at each step and using as $U_{\infty}$ the predicted safety function. Notably, the maximum of the control time series typically aligns closely with the true safety function. However, in regions of minimum control, a lower bound emerges where all orbits exhibit the same maximum control value.
Figure 4: Comparison of predicted safety functions with the true safety function across various models. Each subplot illustrates the similarity of the predictions with their ground truths. As the number of sampled points increases, the predicted safety function aligns more closely with the true safety function. The thin blue line denotes the maximum control for each initial condition over $1000$ iterations, calculated by applying Eq. 4 and using the predicted safety function as $U_{\infty}$.
Figure 5 represents the observed behavior in an orbit with initial condition $x_{n}=0.7$ iterated $1000$ times with the control computed with Eq. 4 at each step. The difference between the two time series is due to the method for estimation of the safety function. In the horizontal axis, the intervals of the safe set appear in pale yellow. The safe set includes all regions in $Q$ where the safety function is less or equal to this value.
Using the true safety function in Fig. 5(a), the safe set identifies regions where the orbit remains stable indefinitely under the suggested control. In Fig. 5(b), the predicted safety function, derived from a model trained with $2000$ points, introduces misalignments in the safe set. Overestimation of the safety function leads to underestimation of the safe set (red regions), while underestimation of the safety function causes overestimation of the safe set (orange regions). A zoomed in view near $x_{n}=0.7$, shown in the top right corner of Fig. 5(b), highlights these discrepancies.
@@
A significant challenge in this model-free approach lies in accurately capturing the lower control regions of the true safety function, particularly when data is sparse. In such scenarios, the estimated safety function often diverges from the true safety function in these regions. This discrepancy arises because sharp variations in the lower control zones depend on data points that are frequently absent during sampling, hindering precise detection and representation. This phenomenon aligns with previous studies on dynamical systems subject to minimal disturbance, which demonstrate that as control decreases, the safe set of orbits within a region $Q$ forms a Cantor-like structure, eventually vanishing from the system [12].
Regarding the performance of the trained models, Fig. 6 illustrates the evolution of the mean squared error (MSE) in estimating safety functions as the noise value $\xi$ in the source maps increases. This analysis involved the computation of $1500$ safety functions derived from random maps. Specifically, for each of the $100$ distinct values of $\xi$ ranging from $0.001$ to $0.1$, $15$ safety functions were generated using maps that shared the same value of $\xi$. For each safety function, the mean MSE was calculated in two steps: first, determining the average MSE across all $1000$ points defining the predicted safety function; and second, averaging these MSE values over the $15$ safety functions.
@@
The results demonstrate that the models estimate safety functions more accurately for maps with lower noise upper bounds, with performance deteriorating as the noise value increases. Nevertheless, the predictions remain reliable across the entire range of noise levels. This trend aligns with a training induced bias, as described in A.2, where systems with lower noise were prioritized to enhance the estimation of the lower control regions of the safety function.
To conclude the analysis of the aforementioned dataset, Table 1 provides a detailed summary of each training model’s performance, including the mean computations time per safety function with $1000$ points and its standard deviation, as well as the mean MSE of the dataset and its standard deviation. For comparison, we also show the computation time of the safety functions obtained through the classical method. As explained in A, these functions are computed using Eq. 11 and evaluated on the same set of points $Q$ as the predicted safety functions. The average computation time is calculated over the $1500$ safety functions of the last dataset and over $50$ iterations of the equation, assuming convergence is reached.
@@
Classical Algorithm $0.75\pm 0.06$
Table 1: Model performance summary: Increasing the number of sampled points generally improves accuracy (lower MSE) but extends computation time. Models with $1000$ or $500$ points achieve similar MSE to the $2000$ point model under evenly distributed noise. Additionally, models with fewer than $500$ points show minimal computation time variation, likely due to hardware constraints discussed in Section 3. They also outperform the classical safety function estimation algorithm in terms of computation time. This table, in conjunction with Fig. 6, highlights the trade-off between model accuracy, prediction time, and the effects of model size on prediction error across different noise levels.
As expected, increasing the amount of information sampled from a system, results in longer prediction times and reduced MSE. However, for models processing $500$ points or fewer, no significant variation in computation time was observed. This consistency in computation time is likely attributed to the hardware specifications used for executing the models, as discussed in Sec. 3. This lower threshold reflects the computational overhead of the programs that can be optimized to meet a application requirement.
@@
5 Conclusion
In this study, we address the fundamental control objective of transforming transient chaos into permanent chaotic motion using only time series. To meet this goal, we propose a transformer-based framework designed to predict the safety functions of dynamical systems without the need for an explicit physical model. This model-free approach offers notable advantages over traditional algorithms. Specifically, our framework achieves competitive accuracy while relying only on the time series of the data. Whereas classical methods require prior knowledge of the system map and set an upper bound of disturbance, our approach, once trained, merely needs a representative time series taken from the unstable dynamical system, to approximate the safety function. Consequently, the control of an orbit depends only on the current perturbed state and the predicted safety function, enabling real-time control.
@@ 5 Conclusion
A key strength of the transformer-based approach is its significantly lower computational complexity compared to classical methods. The classical computation of safety functions, as defined in A, involves an iterative process with a computational complexity of $\mathcal{O}(K\cdot M\cdot N\cdot 2^{d})$, where $K$ is the number of iterations required for the safety function to converge, $M$ is all the possible disturbance scenarios affecting a point, $N$ is the number of neighbors per point, and $d$ is the dimensionality of the system. This iterative nature makes classical methods computationally prohibitive for higher dimensional systems, especially as the number of neighbors or required iterations increases.
@@
We also explored the trade-off between data volume, predictive accuracy, and computation time for safety functions composing of $1000$ points. As shown in Table 1, increasing sample size generally improves accuracy and robustness but at the cost of increased training time and computational demands. These results suggest that selecting an optimal balance between sampled points and computational cost can yield efficient and accurate performance tailored to specific applications. Further research could explore strategies such as adaptive sampling and padding to enhance model performance with limited or imbalanced data.
@@ Appendix A Classical computation of safety functions
This appendix provides a detailed methodology for the classical computation of safety functions. These functions play a critical role in extending the partial control method, enabling the precise determination of minimal control bounds necessary to sustain trajectories within a predefined region $Q$ of the phase space.
@@ A.1 Computation with disturbances
Comments: -Note that the values $u[q_{i},\xi_{s},q_{j}]$ remain constant throughout each iteration of the while loop; therefore, they should be computed once and stored for reuse.
Figure 8: Step-by-step computation of the safety function $U_{k+1}[q_{i}]$ under disturbances. This diagram illustrates how the safety function is updated at a grid point $q_{i}$. (a) A set of disturbances $\xi_{s}$ is applied to $q_{i}$, producing the disturbed images $f(q_{i},\xi_{s})$. (b) From each disturbed point, we compute the control required to reach every other grid point $q_{j}\in Q$. (c) Each control value is compared with the current safety value $U_{k}[q_{j}]$, and for each disturbance, the target $q_{j}$ requiring the least combined effort is selected. (d) Among all disturbances, the worst-case (maximum) of these minimal efforts is used to update the safety function at $q_{i}$ as $U_{k+1}[q_{i}]=\max\big{(}u[q_{i},\xi_{s},q_{j}],U_{k}[q_{j}]\big{)}$.
A.2 Calculation of safety functions
@@ B.1 Calculation of random dynamical systems
The number of functions to be used $M$ is randomly determined from predefined ranges for each set: $4$ to $8$ from Set A, $2$ to $4$ from Set B, $5$ to $8$ from Set C, and $1$ to $2$ from Set D. These ranges are designed to ensure diversity in the generated functions. Once the random functions $g_{i}(x)$ are selected from the respective sets, a weight $c_{i}$ is assigned to each function. These weights are random coefficients uniformly sampled from the range $[0,1]$. The weighted selected functions are then combined into a single expression $f(x)$ as shown in Eq. 12.
$f(x)=\sum_{i=1}^{M}c_{i}g_{i}(x).$ (12)
@@ B.2 Time series simulation
Regarding the system orbits, we iterated $1000$ random initial conditions under the system’s dynamics until they either exited region $Q$ or completed $50$ iterations. The time series evolved according to $x_{n+1}=f(x_{n})+\xi_{n}$. If an orbit exits region $Q$, it is truncated and padded with $-1$ values to achieve a consistent length of $50$ points. This process generates a time series with a shape $(1000,50)$ for each random function.
@@
Set A Set C
$x^{2}(1-x)^{2}$ $\sin^{2}(\pi x)x(1-x)$ $(4+r)x(1-x)^{2}$ $\sqrt{x}(1-x)^{2}$ $x^{3}(1-x)^{3}$ $(1-\cos(\pi x))x(1-x)$ $\tanh^{2}(4x-2)x(1-x)$ $(1-e^{-4x})^{2}x(1-x)$ $(2+r)\min\left(\frac{x}{c},\frac{1-x}{1-c}\right)^{2}x(1-x)$ $\lvert(1-e^{-5x})e^{-5(1-x)}x(1-x)\rvert$ $\lvert\log(1+9x)(1-x)^{2}\rvert$ $\lvert(x-x^{2})^{2}\sin(2\pi x)\rvert$ $\lvert x^{2}(1-x)\cos^{2}(\pi x)\rvert$ $\lvert x(1-x^{4})\rvert$ $\lvert e^{-2(x-0.5)^{2}}x(1-x)\rvert$ $\lvert(\sin(2\pi x)+1)x(1-x)\rvert$ $\lvert(x\log(x+10^{-5}))(1-x)\rvert$ $\lvert\sqrt{x}(1-x^{3})\rvert$ $\lvert x^{3}(1-x^{5})\rvert$
Set B Set D
$\lvert(1-x^{2})\sin^{2}(\pi x)\rvert$ $\lvert(x-x^{2})\cos^{2}(\pi x)\rvert$ $\lvert x(1-x)(1+x-x^{2})\rvert$ $\lvert(1-x)\sin(\pi x)x\rvert$ $\lvert(x(1-x))^{2}\rvert$ $\lvert x^{3}(1-x)(2-x)\rvert$ $\lvert x(1-x)e^{-x}\rvert$ $\lvert(1-x)(1-\cos(2\pi x))\rvert$ $\lvert\sin^{2}(\pi x)x^{2}(1-x)\rvert$ $\lvert(1-e^{-x})(1-x)x\rvert$ $\lvert\cos(\pi x)x^{2}(1-x)\rvert$ $\lvert(x^{2}-x)\cos(\pi x)\rvert$ $\lvert(x(1-x))\sin^{2}(4\pi x)\rvert$ $\lvert(x(1-x))\cos^{2}(6\pi x)\rvert$ $\lvert(x(1-x))(\sin^{2}(2\pi x)+\cos^{2}(2\pi x))\rvert$ $\lvert(x(1-x))e^{-10(x-0.5)^{2}}\rvert$ $\lvert(x(1-x))(1+\cos(10\pi x))\rvert$ $\lvert(x(1-x))\lvert\sin(8\pi x)\rvert\rvert$ $\lvert(x(1-x))(1-\cos^{2}(12\pi x))\rvert$ $\lvert(x(1-x))(1+\sin^{2}(6\pi x))\rvert$ $\lvert(x(1-x))(1+e^{-5(x-0.3)^{2}}+e^{-5(x-0.7)^{2}})\rvert$ $\lvert(x(1-x))(1+\sin(4\pi x)\cos(4\pi x))\rvert$ $\lvert\sin(2\pi x)x(1-x)\rvert$ $\lvert x(1-x)(\text{mod}(10x,1)>0.5)\rvert$ $(3+r)\lvert\text{mod}(10x,1)-0.5\rvert x(1-x)$ $\lvert\lfloor 10x-5\rfloor(1-x)x\rvert$ $(4+r)\lvert\cos(\pi x)\rvert x(1-x)$ $\lvert\lceil 10x-5\rceil(1-x)x\rvert$ $\lvert(\text{mod}(20x,1)>0.5)(1-x)x\rvert$ $\lvert\tan(\pi(x-0.5))(1-x)x\rvert$ $\lvert\text{mod}(\lfloor 5x\rfloor,2)(1-x)x\rvert$ $\lvert\text{sawtooth}(2\pi x)(1-x)x\rvert$ $\lvert\text{sign}(\cos(2\pi x))(1-x)x\rvert$ $\lvert\text{mod}(50x,1)(1-x)x\rvert$ $\lvert\cos(4\pi x)(1-x)x\rvert$
[…]